### v2

# Image Can Bring Your Memory Back: A Novel Multi-Modal Guided Attack against Image Generation Model Unlearning

Abstract
Recent advances in diffusion-based image generation models (IGMs), such as Stable Diffusion (SD), have substantially improved the quality and diversity of AI-generated content. However, these models also pose ethical, legal, and societal risks, including the generation of harmful, misleading, or copyright-infringing material. Machine unlearning (MU) has emerged as a promising mitigation by selectively removing undesirable concepts from pretrained models, yet the robustness of existing methods, particularly under multi-modal adversarial inputs, remains insufficiently explored. To address this gap, we propose Recall, a multi-modal adversarial framework for systematically evaluating and compromising the robustness of unlearned IGMs. Unlike prior approaches that primarily optimize adversarial text prompts, Recall exploits the native multi-modal conditioning of diffusion models by efficiently optimizing adversarial image prompts guided by a single semantically relevant reference image. Extensive experiments across ten state-of-the-art unlearning methods and diverse representative tasks show that Recall consistently surpasses existing baselines in adversarial effectiveness, computational efficiency, and semantic fidelity to the original prompt. These results reveal critical vulnerabilities in current unlearning pipelines and underscore the need for more robust, verifiable unlearning mechanisms. More than just an attack, Recall also serves as an auditing tool for model owners and unlearning practitioners, enabling systematic robustness evaluation. Code and data are available at https://github.com/ryliu68/RECALL.
Warning: This paper contains visual content that may include explicit or sensitive material, which some readers may find disturbing or offensive.
1 Introduction
The emergence of image generation models (IGMs), such as Stable Diffusion (Rombach et al., 2022a), has greatly advanced the quality and diversity of AI-generated visual content. IGMs are now widely used in digital art, multimedia creation, and visual storytelling (Chen et al., 2024; Zhang et al., 2024b). However, their rapid adoption also raises serious ethical and legal concerns, particularly regarding the misuse of these models to generate harmful, misleading, or infringing content (Qu et al., 2023; Schramowski et al., 2023). Consequently, ensuring robust safety and trustworthiness mechanisms within these generative frameworks has emerged as an urgent imperative (Wang et al., 2025; 2024).
Among different lines of efforts, machine unlearning (MU) has recently gained growing prominence (Zhang et al., 2024c; Park et al., 2024; Li et al., 2024b). It aims to remove sensitive concepts (e.g., nudity, violence, and copyrighted materials) from the IGMs, prohibiting the generation of sensitive or problematic content while maintaining the model’s general capability of producing benign and high-quality outputs (Schramowski et al., 2023; Kumari et al., 2023; Gandikota et al., 2024). Recent IGM unlearning (IGMU) methods utilize diverse strategies, including fine-tuning (Gandikota et al., 2023; Zhang et al., 2024a), targeted concept removal (Gandikota et al., 2024; Gong et al., 2024; Orgad et al., 2023), negative prompting (Schramowski et al., 2023), and adversarial filtering (Zhang et al., 2024c; Gong et al., 2024; Wu et al., 2025). They have proven effective in safety protection of contemporary IGMs, enforcing compliance with ethical guidelines and legal standards.
Despite the rapid progress in this field, the practical robustness of these techniques is challenged, especially under adversarial scenarios. Recent studies have revealed that unlearned IGMs are still vulnerable: carefully optimized prompts can successfully circumvent safety mechanisms, prompting the unlearned models to regenerate prohibited content (Zhang et al., 2024d; Tsai et al., 2024). However, these attack methods mainly focus on perturbing the textual modality and suffer from the following critical limitations. ① Modifying textual inputs can disrupt the semantic alignment between the generated images and original prompts; ② Many approaches rely on external classifiers or additional diffusion models for adversarial text prompt optimization, incurring substantial computational overhead; ③ Their effectiveness sharply declines against robust, adversarially-enhanced unlearning methods, e.g., AdvUnlearn (Zhang et al., 2024c), RECE (Gong et al., 2024); ④ Crucially, these methods overlook the inherent multi-modal conditioning capabilities (e.g., simultaneous textual and image) of IGMs, thus missing a critical dimension of potential vulnerability.
Figure 1: Given an assumed successfully unlearned IGM $\mathcal{G}_{u}$, our adversarial image prompt $P^{adv}_{img}$ combined with the original sensitive text prompt $P_{text}$ as multi-modal guidance can circumvent the unlearning mechanism, leading to the reappearance of removed content $I^{*}$. Sensitive parts are covered by Refer to caption.
To address these limitations, we propose Recall, a novel multi-modal attack framework against mainstream IGMU solutions. Figure 1 illustrates the attack scenarios. First, unlike previous attacks that focus solely on text perturbation, Recall strategically integrates an adversarially optimized image with the original text prompt to attack the unlearned model, ensuring strong semantic alignment between the generated images and corresponding textual descriptions. Second, Recall performs the attack within the unlearned model and optimizes the latent representation of the adversarial image prompt, eliminating the reliance on additional components and significantly enhancing computational efficiency. Furthermore, by introducing adversarial perturbations directly within the image modality, Recall effectively exposes hidden vulnerabilities in adversarially enhanced unlearning methods, revealing their susceptibility to image-based attacks that prior text-based adversarial techniques may overlook. Finally, Recall fully exploits the inherent multi-modal guidance capabilities of IGMs, enabling the comprehensive identification of critical vulnerabilities across diverse scenarios before real-world deployment.
Extensive empirical results conducted on ten state-of-the-art IGMU methods across four representative unlearning scenarios demonstrate that Recall consistently surpasses prior approaches in terms of adversarial effectiveness, computational efficiency, and semantic fidelity. Beyond demonstrating strong attack performance, these findings reveal critical vulnerabilities in current unlearning pipelines, underscoring their susceptibility to multi-modal guided adversarial inputs and the urgent need for more robust and verifiable unlearning mechanisms in IGMs. From the perspective of model owners, Recall can also serve as an efficient robustness auditing tool to assess the effectiveness of their unlearning procedures. Our key contributions are as follows:
  • We propose Recall, the first multi-modal guided attack framework to break the robustness of IGMU techniques, allowing the protected model to regenerate unlearned sensitive concepts with high semantic fidelity.
  • Recall introduces a highly efficient optimization strategy that operates solely within the unlearned model by utilizing only a single reference image, eliminating the need for auxiliary classifiers, original diffusion models, or external semantic guidance required by previous attacks.
  • Through comprehensive experiments covering ten representative IGMU techniques across four diverse tasks, we empirically demonstrate the vulnerabilities of existing unlearning solutions under multi-modal attacks, revealing the urgent need for more robust safety unlearning.
2 Related Work
Image Generation Models (IGMs). Diffusion-based IGMs, such as Stable Diffusion (SD) (Rombach et al., 2022b), DALL·E (OpenAI, 2023a), and Imagen (Saharia et al., 2022), have achieved impressive progress in synthesizing diverse, high-fidelity images. These models leverage large-scale datasets (e.g., LAION-5B (Schuhmann et al., 2022)) and integrate components including pre-trained text encoders (e.g., CLIP (Radford et al., 2021)), U-Net denoisers, and VAE decoders, enabling precise semantic alignment between prompts and images for a wide range of applications.
Unlearning in Image Generation Models. The proliferation of IGMs has led to increasing concerns about the generation of harmful or copyrighted content (Qu et al., 2023; Liu et al., 2025). Machine unlearning (MU) methods have been developed to selectively remove undesirable concepts from pretrained models (Kumari et al., 2023; Fan et al., 2024) while preserving overall generative capabilities. Existing IGM unlearning (IGMU) approaches can be broadly categorized as: (i) Fine-tuning-based, which update model parameters to forget specific concepts (e.g., ESD (Gandikota et al., 2023), UCE (Gandikota et al., 2024)); (ii) Guidance-based, which constrain generation at inference without modifying model weights (e.g., SLD (Schramowski et al., 2023)); and (iii) Regularization-based, which introduce forgetting objectives during training (e.g., Receler (Huang et al., 2024), FMN (Zhang et al., 2024a)). Despite their successes, these methods often exhibit limited robustness and generalization.
Adversarial Attacks on IGMU. Recent works demonstrate that adversarial prompts can circumvent IGMU defenses and recover restricted content (Li et al., 2024a; Chin et al., 2024a; Zhang et al., 2024d). White-box methods such as P4D (Chin et al., 2024a) and UnlearnDiffAtk (Zhang et al., 2024d) optimize text prompts but may suffer from high computational overhead and reduced semantic alignment, CCE (Pham et al., 2024) learns a single placeholder via textual inversion on the erased model and substitutes it at inference to recover restricted concepts, while WACE (Lu et al., 2025) regenerates forgotten content via a noise-based probe. The black-box and transfer-based approaches (Han et al., 2025; Tsai et al., 2024; Dang et al., 2025; Ma et al., 2025; Chin et al., 2024b) rely on surrogate models or textual perturbations, sometimes requiring external classifiers or access to the original diffusion model. These approaches are less effective against stronger unlearning defenses (e.g., AdvUnlearn (Zhang et al., 2024c), Receler (Huang et al., 2024)) and remain computationally intensive.
Therefore, effective attack strategies should efficiently recover restricted content, preserve prompt-image semantic coherence, and exploit vulnerabilities beyond text perturbations. To this end, we propose Recall, a multi-modal adversarial framework that leverages adversarial image prompts alongside unmodified text inputs, enabling effective attacks on unlearned models via multi-modal guidance. Our method requires neither external classifiers nor access to the original IGM, making it both lightweight and effective.
3 Preliminary
3.1 Image Generation Model Unlearning
Given a pretrained IGM $\mathcal{G}$ over a concept space $\mathcal{C}$, Image Generation Model Unlearning (IGMU) aims to selectively remove the model’s ability to generate content associated with a sensitive concept subset $\mathcal{C}^{\prime}\subseteq\mathcal{C}$, while preserving generative quality for the remaining concepts. Formally, an unlearning algorithm $\mathcal{A}_{u}$ produces a modified model $\mathcal{G}_{u}=\mathcal{A}_{u}(\mathcal{G},\mathcal{C}^{\prime})$. The unlearning objectives are twofold:
  • Forgetting: For all $c\in\mathcal{C}^{\prime}$, the model should no longer generate content related to $c$: $\mathcal{G}_{u}(P_{text})\cap\mathcal{G}(P_{text})=\emptyset$.
  • Preservation: For all $c\in\mathcal{C}\setminus\mathcal{C}^{\prime}$, the generative performance should be retained: $sim\bigl(\mathcal{G}_{u}(P_{text}),\mathcal{G}(P_{text})\bigr)\geq\sigma$, where $sim(\cdot,\cdot)$ denotes a perceptual similarity metric (e.g., CLIP score or LPIPS), and $\sigma$ is a predefined threshold.
In this work, we focus on unlearning methods and evaluation within the multi-modal, diffusion-based IGM setting, with SD as a representative backbone.
3.2 Threat Model
We consider an adversary seeking to deliberately regenerate erased content from a concept-unlearned, multi-modal (text+image) IGM. The adversary requires white-box access and the ability to invoke the model’s native multi-modal-conditioning pathway. This setting primarily targets (i) attacks: it is realistic because many applications deploy open or publicly available Stable Diffusion variants and is consistent with prior white-box threat models (Chin et al., 2024a; Pham et al., 2024; Zhang et al., 2024d). The same setup also supports (ii) unlearning red-teaming by model owners or auditors as a pre-deployment robustness assessment to locate weaknesses and guide verifiable mitigation.
3.3 Problem Formulation
We introduce a new attack strategy that optimizes image prompts by leveraging multi-modal guidance, which is natively supported by Stable Diffusion (Rombach et al., 2022a), to bypass unlearning mechanisms and regenerate erased content.
Given an unlearned image generation model (IGM) $\mathcal{G}_{u}$ that has been updated to suppress content associated with target concept $c$, a text prompt $P_{text}$ containing $c$, and an image $P_{img}$ relevant to the concept $c$, we aim to find an adversarial image input $P_{img}^{adv}$ such that, when paired with $P_{text}$, the unlearned IGM $\mathcal{G}_{u}$ generates image $I^{*}$ related to $c$:
$I^{*}=\mathcal{G}_{u}(P_{img}^{adv},P_{text}),\quad\text{s.t.}\quad I^{*}\approx I\mid c,$ (1)
where $I\mid c$ denotes images that explicitly contains the target concept $c$; these images can come from the original model $\mathcal{G}$ with the same text prompt $P_{\text{text}}$ or from any other source.
The adversarial image prompt $P_{img}^{adv}$ is obtained by solving:
$P_{img}^{adv}=\arg\min_{P_{img}}\mathcal{L}_{adv}\left(\mathcal{G}_{u}(P_{img},P_{text}),I\right),$ (2)
where $\mathcal{L}_{adv}$ is an adversarial loss function.
Unlike prior attacks that modify the text prompt $P_{text}$, we optimize $P_{img}$ while keeping $P_{text}$ unchanged, thus preserving the semantic intent. The optimization follows a gradient-based approach:
$P_{img}^{adv}\leftarrow P_{img}-\eta\cdot\nabla_{P_{img}}\mathcal{L}_{adv}(\mathcal{G}_{u}(P_{img},P_{text}),I),$ (3)
where $\eta$ is the step size. This process enables the adversarial image prompt $P_{img}^{adv}$, together with $P_{text}$, to exploit vulnerabilities in the unlearned model and recover the erased content while maintaining semantic alignment with the text prompt.
4 Methodology
4.1 Overview
We propose Recall, a multi-modal adversarial framework targeting unlearned IGMs. Unlike conventional text-only attacks, Recall jointly optimizes adversarial image prompt by leveraging a reference image $P_{ref}$ as guidance. As illustrated in Figure 2, Recall comprises three stages: (1) Latent Encoding: The reference image $P_{ref}$ and a noise-injected initial prompt are encoded into latent representations. (2) Iterative Latent Optimization: The adversarial latent is iteratively refined under the guidance of the reference latent by minimizing the discrepancy between their predicted noise residuals. (3) Multi-modal Attack: The optimized latent is decoded to an adversarial image, which, paired with the text prompt, forms a multi-modal input to the unlearned IGM, enabling effective recovery of the erased target concept. Details of each stage are described in the following sections, and the overall pipeline is summarized in Algorithm 1 (Appendix B).
Figure 2: Overview of Recall. Given a reference image $P_{ref}$ that depicts the erased concept and a heavily noised initial image prompt $P_{img}^{init}$, we iteratively optimize the latent $z_{adv}$ (initialized from $P_{img}^{init}$) to align with the reference latent $z_{ref}$ under the same text condition. After optimization, $z_{adv}$ is decoded into an adversarial image $P_{img}^{adv}$, which is then paired with the original text prompt and fed into the unlearned model, enabling recovery of the erased concept, thereby exposing vulnerabilities of current unlearning mechanisms under multi-modal guidance.
4.2 Image Encoding
To avoid incurring additional computational overhead from external classifiers or relying on the original IGM, we introduce a reference image $P_{ref}$ containing the target concept $c$ to guide the generation process, where the reference image $P_{ref}$ can be sourced from multiple sources, such the internet, public datasets, or self-collected. This reference implicitly embeds the erased concept, thereby facilitating adversarial optimization of the initial image prompt $P_{img}^{init}$. To enhance efficiency and precision, Recall performs the optimization directly in the latent space representation $z_{adv}$ of the image prompt.
As illustrated in Figure 2, we initialize $P_{img}^{init}$ by blending a small portion of the reference image $P_{ref}$ with random noise $\delta$ sampled from an isotropic Gaussian distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$:
$P_{img}^{init}\leftarrow\lambda\cdot P_{ref}+(1-\lambda)\cdot\delta,\quad\delta\sim\mathcal{N}(0,I),$ (4)
where $\lambda\in[0,1]$ is a hyperparameter controlling the semantic similarity to the reference image. We set $\lambda=0.25$ throughout our experiments. This approach increases the sampling space of Stable Diffusion and further enhances the diversity of the generated images, while simultaneously encouraging the generation process to better follow the guidance of the text prompt, thereby improving semantic consistency.
To accelerate optimization, both $P_{img}^{init}$ and $P_{ref}$ are encoded into the latent space using the image encoder $\mathcal{E}_{i}$ from the unlearned model, yielding:
$z_{i}=\mathcal{E}_{i}(P_{img}^{init}),\quad z_{ref}=\mathcal{E}_{i}(P_{ref}),$ (5)
where $z_{i}$ is used as the initial adversarial latent $z_{adv}$, and $z_{ref}$ serves as the fixed reference guiding the optimization process.
4.3 Iterative Latent Optimization
We iteratively optimize the adversarial latent as below.
Generation of Latent $z_{t}$.
Unlike standard latent diffusion, which typically initializes from a randomly sampled latent, Recall generates the noisy latent at timestep $t$ as:
$z_{t}=\sqrt{\bar{\alpha}_{t}}\,z+\sqrt{1-\bar{\alpha}_{t}}\,\epsilon,\quad\epsilon\sim\mathcal{N}(0,I),$ (6)
where $z$ denotes either the reference latent $z_{ref}$ or the adversarial latent $z_{adv}$. The cumulative noise schedule $\bar{\alpha}_{t}$ determines the relative contribution of signal and noise.
To accelerate optimization, each $z_{t}$ corresponds to a single denoising step from a fixed DDIM (Song et al., 2021) sampling schedule of 50 steps ($t=T\rightarrow 0$). At each step, we apply one backward denoising pass to simulate efficient adversarial guidance. We adopt an early stopping mechanism: the attack halts as soon as the target content reappears; It fails if no target content is observed after all steps are exhausted.
Optimization under Multi-Modal Guidance.
For each noisy latent $z_{t}$, the diffusion model predicts the corresponding noise component using a U-Net $\mathcal{F}_{\theta}$, conditioned on the textual embedding $h_{t}$ from the encoding text prompt $P_{text}$ by the text encoder $\mathcal{E}_{t}$ (i.e., $h_{t}=\mathcal{E}_{t}(P_{text})$). The predicted noise of reference image $\hat{\epsilon}_{ref}$ and adversarial image $\hat{\epsilon}_{adv}$ can be derived as:
$\hat{\epsilon}_{ref}=\mathcal{F}_{\theta}(z_{\{ref,t\}},t,h_{t});\ \hat{\epsilon}_{adv}=\mathcal{F}_{\theta}(z_{\{adv,t\}},t,h_{t}).$ (7)
The discrepancy between two noise predictions forms the basis of the adversarial objective function.
As discussed previously, our attack explicitly targets the latent representation $z_{adv}$ of the adversarial image prompt $P_{img}^{adv}$, aiming to efficiently induce the unlearned IGM model to regenerate the previously unlearned content. Specifically, at each diffusion timestep $t$, we iteratively refine the adversarial latent representation $z_{adv}$ using a gradient-based optimization procedure guided by the adversarial loss $\mathcal{L}_{adv}$. To enhance stability and facilitate convergence, we incorporate momentum-based gradient normalization into our optimization scheme (Dong et al., 2018). Specifically, we iteratively update the latent adversarial variable $z_{adv}$ over $N$ epochs according to:
$v_{i}=\beta\cdot v_{i-1}+\frac{\nabla_{z_{adv}}\mathcal{L}_{adv}}{\|\nabla_{z_{adv}}\mathcal{L}_{adv}\|_{1}+\omega},\ z_{adv}\leftarrow z_{adv}+\eta\cdot\text{sign}(v_{i}),$ (8)
where $\eta$ denotes the step size, $v_{i}$ is the momentum-updated gradient direction at iteration $i$, and $\beta=0.9$ represents the momentum factor. The term $\nabla_{z_{adv}}\mathcal{L}_{adv}$ refers to the gradient of the adversarial loss $\mathcal{L}_{adv}$ with respect to the adversarial latent $z_{adv}$, normalized by its $L_{1}$-norm for gradient scale invariance, and $\omega=1\mathrm{e}{-8}$ is a small constant for numerical stability. Furthermore, in practical implementations, we periodically integrate a small portion of the reference latent $z_{ref}$ back into $z_{adv}$, thereby reinforcing semantic consistency between $z_{adv}$ and $z_{ref}$ during the optimization:
$z_{adv}\leftarrow(1-\gamma)z_{adv}+\gamma\cdot z_{ref},$ (9)
where $\gamma$ is a small regularization parameter and set to 0.05 in our optimization.
Objective Function.
The adversarial objective function $\mathcal{L}_{adv}$ explicitly quantifies the discrepancy between noise predictions generated from the adversarial latent $\hat{\epsilon}_{adv}$ and reference latent $\hat{\epsilon}_{ref}$ with U-Net at step $t$, respectively:
[…]
where $\mathcal{M}$ denotes a similarity measurement. In this work, we employ the mean squared error (MSE).
Adversarial Image Reconstruction.
After optimization, the refined adversarial latent $z_{adv}$ is subsequently decoded into the image space through the image decoder $\mathcal{D}_{i}$ of the unlearned SD model to generate the final adversarial image used for the attack: $P_{img}^{adv}=\mathcal{D}_{i}(z_{adv})$.
4.4 Multi-modal Attack
Once the adversarial image $P_{img}^{adv}$ is obtained, we leverage the multi-modal conditioning mechanism of the unlearned model $\mathcal{G}_{u}$ to generate images containing the forgotten content and semantically aligned with the text prompt $P_{text}$. The final image generation process integrates both the optimized adversarial image prompt and the original text prompt in a multi-modal manner:
$I^{*}=\mathcal{G}_{u}(P_{img}^{adv},P_{text}),$ (11)
where $I^{*}$ is the final generated image.
Our method systematically exposes the inherent weaknesses in current concept unlearning techniques: by utilizing both adversarial image optimization and textual conditioning, the unlearned information can still be reconstructed.
Table 1: Attack comparisons against unlearned IGMs in six dataset for four representative unlearning tasks.
Task Method ESD FMN SPM AdvUnlearn MACE RECE DoCo UCE Receler ConceptPrune Avg. ASR
Nudity-I2P Text-only 10.56 66.90 32.39 1.41 3.52 7.04 30.99 8.45 8.45 73.24 24.30
Image-only 0.00 18.31 12.68 4.23 5.63 14.08 3.52 11.97 6.34 13.38 9.01
Text & R_noise 0.70 29.58 14.08 0.70 3.52 1.41 14.79 2.82 0.70 36.62 10.49
Text & Image 13.38 59.15 42.25 7.04 10.56 14.79 40.14 17.61 20.42 52.11 27.74
P4D-K 51.41 80.28 76.76 6.34 40.14 35.92 77.46 56.34 40.14 77.46 54.22
P4D-N 62.68 88.73 76.76 2.82 32.39 52.11 80.28 54.93 35.92 89.44 57.61
CCE 59.15 85.21 64.08 37.32 57.75 26.76 30.28 40.14 20.42 83.10 50.42
UnlearnDiffAtk 51.41 92.25 88.03 8.45 47.18 40.85 87.32 70.42 55.63 97.18 63.87
WACE-N 30.28 80.99 61.27 4.23 20.42 15.49 58.45 28.17 23.24 80.28 40.28
WACE-C 51.41 89.44 79.58 25.35 46.48 28.87 71.83 42.96 46.48 88.03 57.04
Recall 71.83 100.00 96.48 60.56 71.83 59.86 92.25 76.76 78.87 99.30 80.77
Nudity-MMA Text-only 1.56 46.88 32.03 0.00 0.00 13.28 27.34 24.22 14.06 53.12 21.25
Text & R_noise 0.00 20.31 17.19 0.00 0.00 4.69 21.88 14.84 1.56 42.19 12.27
Text & Image 8.59 78.91 59.38 0.00 2.34 37.50 62.50 44.53 43.75 80.47 41.80
P4D-K 56.90 62.50 76.88 8.43 49.54 37.50 85.31 80.62 79.69 89.65 62.70
P4D-N 62.50 74.37 78.44 10.64 53.67 51.25 88.44 91.41 85.94 98.44 69.51
CCE 35.16 89.84 78.91 3.12 55.47 46.88 54.69 58.59 36.72 97.66 55.70
UnlearnDiffAtk 40.62 100.00 99.22 23.78 33.59 89.06 98.44 95.31 85.94 99.22 76.52
WACE-N 28.12 86.72 75.78 7.03 1.56 46.09 68.75 49.22 52.34 86.72 50.23
WACE-C 61.72 92.19 82.81 49.22 13.28 57.03 80.47 70.31 70.31 85.16 66.25
Recall 75.78 100.00 97.66 82.81 53.12 89.84 94.53 92.97 96.09 99.22 88.20
Nudity-ART Text-only 0.00 11.72 2.34 0.00 0.00 0.00 3.12 0.78 2.34 7.03 2.73
Text & R_noise 0.00 35.16 17.19 0.00 0.00 0.00 4.69 0.78 0.00 17.19 7.50
Text & Image 0.78 14.84 13.28 0.78 1.56 1.56 4.69 2.34 4.69 12.50 5.70
P4D-K 8.73 66.67 57.63 2.86 31.45 28.49 43.87 41.83 12.50 56.25 35.03
P4D-N 12.50 62.86 62.81 3.98 24.69 32.81 48.44 45.62 21.88 53.12 36.87
CCE 20.31 53.91 28.12 28.12 21.88 3.12 3.12 13.18 6.25 42.97 22.10
UnlearnDiffAtk 31.25 76.56 66.41 0.78 17.19 21.09 76.47 39.06 35.16 75.78 43.98
WACE-N 7.81 48.44 26.56 0.78 2.34 4.69 21.09 7.03 7.81 37.50 16.41
WACE-C 20.31 57.81 34.38 5.47 8.59 4.69 33.59 10.94 14.84 47.66 23.83
Recall 62.50 91.29 81.25 32.03 43.75 32.99 98.12 52.34 70.31 89.84 65.44
Van Gogh-style Text-only 26.00 50.00 82.00 24.00 72.00 74.00 52.00 98.00 20.00 98.00 59.60
Image-only 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Text & R_noise 8.00 14.00 18.00 12.00 16.00 28.00 38.00 38.00 10.00 80.00 26.20
Text & Image 10.00 18.00 42.00 10.00 24.00 32.00 42.00 74.00 24.00 96.00 37.20
P4D-K 56.00 72.00 90.00 86.00 82.00 100.00 62.00 94.00 62.00 98.00 80.20
P4D-N 88.00 88.00 100.00 86.00 96.00 98.00 90.00 100.00 74.00 100.00 92.00
CCE 78.00 66.00 100.00 86.00 94.00 88.00 46.00 98.00 16.00 100.00 77.20
UnlearnDiffAtk 96.00 100.00 100.00 84.00 100.00 100.00 100.00 100.00 92.00 100.00 97.20
WACE-N 28.00 36.00 80.00 12.00 70.00 58.00 80.00 90.00 10.00 96.00 56.00
WACE-C 14.00 34.00 86.00 8.00 28.00 50.00 72.00 88.00 4.00 96.00 48.00
Recall 92.00 100.00 100.00 92.00 100.00 100.00 98.00 100.00 92.00 100.00 97.40
Object-Church Text-only 16.00 52.00 44.00 0.00 4.00 4.00 44.00 6.00 2.00 92.00 26.40
Image-only 4.00 18.00 20.00 8.00 16.00 18.00 12.00 20.00 16.00 20.00 15.20
Text & R_noise 0.00 32.00 22.00 0.00 0.00 2.00 32.00 2.00 0.00 46.00 13.60
Text & Image 46.00 66.00 66.00 4.00 10.00 4.00 60.00 8.00 2.00 80.00 34.60
P4D-K 6.00 56.00 48.00 0.00 2.00 28.00 86.00 24.00 20.00 88.00 35.80
P4D-N 58.00 90.00 86.00 14.00 48.00 12.00 92.00 10.00 14.00 74.00 49.80
CCE 54.00 92.00 76.00 58.00 60.00 12.00 58.00 46.00 26.00 76.00 55.80
UnlearnDiffAtk 70.00 96.00 94.00 4.00 32.00 52.00 100.00 66.00 10.00 100.00 62.40
WACE-N 48.00 66.00 60.00 2.00 4.00 6.00 68.00 6.00 2.00 74.00 33.60
WACE-C 58.00 76.00 74.00 8.00 6.00 10.00 80.00 16.00 0.00 86.00 41.40
Recall 96.00 100.00 98.00 62.00 50.00 46.00 98.00 68.00 20.00 98.00 73.40
Object-Parachute Text-only 4.00 54.00 24.00 4.00 2.00 2.00 8.00 2.00 2.00 88.00 19.00
Image-only 20.00 92.00 96.00 88.00 92.00 86.00 96.00 90.00 88.00 84.00 83.20
Text & R_noise 4.00 48.00 22.00 2.00 4.00 0.00 10.00 2.00 2.00 60.00 15.40
Text & Image 94.00 98.00 88.00 52.00 72.00 48.00 50.00 60.00 32.00 98.00 69.20
P4D-K 6.00 40.00 24.00 2.00 4.00 14.00 72.00 18.00 20.00 96.00 29.60
P4D-N 36.00 82.00 70.00 8.00 22.00 12.00 52.00 14.00 2.00 84.00 38.20
CCE 74.00 92.00 72.00 48.00 54.00 34.00 52.00 52.00 38.00 88.00 60.40
UnlearnDiffAtk 56.00 100.00 94.00 14.00 36.00 34.00 92.00 42.00 30.00 100.00 59.80
WACE-N 30.00 84.00 46.00 10.00 10.00 6.00 26.00 4.00 6.00 88.00 31.00
WACE-C 56.00 84.00 60.00 6.00 16.00 8.00 32.00 22.00 14.00 90.00 38.80
Recall 100.00 100.00 100.00 94.00 100.00 88.00 98.00 96.00 94.00 100.00 97.00
5 Experiments
We conduct extensive experiments involving TEN SOTA unlearning techniques across four representative unlearning tasks: Nudity, Van Gogh-style, Object-Church, and Object-Parachute, thus yield a total of forty unlearned IGMs. Our objective is to systematically validate the effectiveness and generalization of our proposed multi-modal guided attack Recall against different scenarios.
5.1 Experimental Setup
Datasets. We evaluate on three nudity unlearning datasets (I2P (Schramowski et al., 2023), MMA (Yang et al., 2024), and ART (Li et al., 2024a)). For the remaining targets, such as Van Gogh-style, Object-Church, and Object-Parachute, we reuse the text prompts released by UnlearnDiffAtk to ensure protocol comparability. For methods that require a reference image, we provide one same additional image per unlearning task. Details of all prompts and reference images are provided in Appendix C.1, Table 3.
IGMU Methods. We evaluate our approach across ten state-of-the-art IGMU techniques: ESD (Gandikota et al., 2023), FMN (Zhang et al., 2024a), SPM (Lyu et al., 2024), AdvUnlearn (Zhang et al., 2024c), MACE (Lu et al., 2024), RECE (Gong et al., 2024), DoCo (Wu et al., 2025), Receler (Huang et al., 2024), ConceptPrune (Chavhan et al., 2025), and UCE (Gandikota et al., 2024). The weights of involved unlearned SD models are sourced from three primary origins: ① the AdvUnlearn GitHub repository [^1] https://github.com/OPTML-Group/AdvUnlearn, as described in (Zhang et al., 2024c); ② weights officially released by their respective authors, such as RECE (Gong et al., 2024), MACE (Lu et al., 2024) and DoCo (Wu et al., 2025); and ③ weights trained in-house using official implementations provided by ourselves.
Baselines. We compare our proposed Recall against several representative attack baselines: Text-only (text prompts only), Image-only (reference image prompt only), Text & R_noise (text with a noised image), Text & Image (text prompt and reference image), P4D (with two variants P4D-K and P4D-N) (Chin et al., 2024a), CCE (Pham et al., 2024), UnlearnDiffAtk (Zhang et al., 2024d), and WACE(with two variants WACE-N and WACE-C) (Lu et al., 2025). Their detailed descriptions and implementation can be found in Appendix C.2.
Evaluation Metrics. We assess the effectiveness of our attack using task-specific deep learning-based detectors and classifiers, including the NudeNet detector (Praneeth, 2023), a ViT-based style classifier (Zhang et al., 2024d), and an ImageNet-pretrained ResNet-50 (He et al., 2016). The primary metric is attack Success Rate (ASR, %) and average ASR for attack performance, average attack time (seconds, s) for computational efficiency, CLIP Score (Hessel et al., 2021) for quantifying semantic alignment between generated images and prompts, and LPIPS (Zhang et al., 2018), Inception Score (IS) (Salimans et al., 2016), and a DINO-based feature distance (Oquab et al., 2024) for generated image diversity. Throughout all tables, the best attack performance is highlighted in bold, while the second-best is indicated with underlining.
[…]
5.2 Attack Performance
Figure 3: Generated images under different attacks. Rows (top to bottom): Nudity, Van Gogh, Church, and Parachute.
We comprehensively evaluate the effectiveness of Recall against several baseline attack methods across four representative unlearning tasks. The detailed experimental results, as summarized in Table 1, reveal several critical findings. ① Existing unlearning approaches fail to fully erase target concepts; notably, original textual or combined text-image prompts (reference image or randomly initialized) alone achieve substantial ASRs. For instance, combined text-image prompts yield an Avg. ASR exceeds 69.20% in the Parachute task. ② All baseline attack methods exhibit limited effectiveness when attacking adversarially enhanced unlearning strategies (e.g., AdvUnlearn and RECE), evidenced by their significantly lower ASRs. ③ In contrast, Recall consistently attains superior performance, achieving average ASRs ranging from 73.40% to 97.40% across diverse scenarios. Specifically, Recall outperforms UnlearnDiffAtk, a strong baseline, improving the average ASR by 16.90%, 0.20%, 11.00%, and 37.20% for four tasks. These results highlight the robustness and efficacy of Recall in regenerating targeted, presumably erased visual concepts.
In addition, qualitative generation results on MACE in Figure 3 (Complete results in Appendix D Table 4) and visual cases (in Appendix F.2 Figure 7) show that Recall consistently surpasses existing baselines in recovering erased concepts across a variety of unlearning scenarios, yielding highly diverse outputs.
5.3 Attack Efficiency
Figure 4: Comparison of average attack time for different attack methods for Nudity task.
To assess the practical efficiency of Recall, we compare the average attack time (in seconds) needed by Recall with various baselines, a lower average attack time indicates higher efficiency. Figure 4 reports results for Nudity task (attack time for more tasks can be found in Appendix E). As shown, Recall achieves significantly lower attack time ($\sim$64s) compared to P4D-N ($\sim$238s), UnlearnDiffAtk ($\sim$232s) and WACE-C ($\sim$243.15s). This improvement stems from our efficient multi-modal optimization directly in the latent space and these efficiency gains align with our high attack success rates, highlighting that Recall is both effective and computationally lightweight. Notably, less robust unlearning methods (e.g., FMN, SPM) tend to require shorter attack durations, further illustrating their susceptibility.
5.4 Semantic Alignment
Table 2: CLIP Scores ($\uparrow$) among baselines and Recall.
Task Method ESD MACE RECE UCE Receler DoCo
Nudity-I2P P4D 24.09 23.20 24.99 24.90 25.64 23.70
CCE 19.11 19.00 19.07 19.08 19.02 19.11
UnlearnDiffAtk 29.61 23.11 29.25 29.17 29.00 31.18
WACE 19.19 18.77 19.83 19.47 19.61 19.25
Recall 32.13 24.79 30.66 31.31 31.12 31.95
Van Gogh P4D 17.73 31.66 25.64 22.57 13.49 21.81
CCE 18.39 18.54 18.43 18.36 18.58 18.40
UnlearnDiffAtk 29.23 33.85 33.10 33.32 21.26 22.39
WACE 18.21 18.04 18.18 18.27 17.98 18.29
Recall 35.92 35.28 34.71 34.20 23.37 30.01
Church P4D 25.88 28.44 27.68 27.76 30.34 25.62
CCE 20.46 20.46 20.50 20.52 20.34 20.64
UnlearnDiffAtk 27.68 27.46 27.04 28.97 30.89 29.99
WACE 20.09 21.16 19.40 19.88 20.53
Recall 27.94 28.94 28.36 27.82 27.73 30.37
Parachute P4D 23.50 23.73 28.01 27.13 24.18 28.37
CCE 21.92 22.03 22.09 22.03 22.03 22.06
UnlearnDiffAtk 25.64 25.59 25.73 23.37 26.22 28.98
WACE 22.05 21.84 21.64 21.97 21.95 21.76
Recall 29.64 28.66 31.04 31.10 28.92 30.63
The original SD model achieves CLIP Scores of 31.05, 33.98, 30.75, and 31.56 on the four tasks, respectively.
We assess the semantic consistency between regenerated images and their corresponding text prompts using the CLIP Score. Table 2 presents the average CLIP Scores for four attack methods, P4D, CCE, UnlearnDiffAtk, WACE, and our proposed Recall, evaluated across six unlearning techniques and four aforementioned representative unlearning tasks.
As shown in Table 2, Recall consistently outperforms baseline methods, achieving the highest CLIP Scores across all tasks and unlearning settings. Notably, Recall attains an average CLIP Score of 30.28, surpassing UnlearnDiffAtk (28.00), P4D (25.00), CCE (20.01) and WACE (19.89). These results indicate that text-based methods, which perturb original prompts, often degrade semantic coherence. In contrast, our multi-modal adversarial framework preserves the textual intent and introduces perturbations solely through the image modality, yielding superior semantic alignment.
5.5 Generalizability
Reference Independence. We assess the robustness of Recall to reference image selection using three additional references ($R_{1}$$R_{3}$, see Appendix F.1, Figure 6). These additional references are randomly downloaded from the Internet for the Nudity and Object-Church tasks; $R_{\text{org}}$ is the default reference in our core experiments, while $R_{1}$$R_{3}$ are used only for this ablation. As reported in Appendix F.1, Table 5, both attack and diversity metrics remain consistently high provided the reference is representative, demonstrating that Recall does not depend on any specific image.
Generation Diversity. We quantitatively compare image-only, text-only, and Recall. As detailed in Appendix F.2, Recall achieves substantially greater diversity than image-only baselines and matches the performance of text-only approaches, indicating that it recovers the original concept distribution rather than simply transforming reference images.
Model Version Independence. We further evaluate Recall on unlearned models based on SD 2.0 and SD 2.1, in addition to SD 1.4. As shown in Appendix F.3, Recall consistently maintains high effectiveness across all versions, confirming its robustness and generalizability to more advanced diffusion architectures.
5.6 Ablation Study
We conduct ablation studies to systematically evaluate the impact of key strategies and hyperparameters on the performance of the Recall framework.
Strategies. We analyze three core strategies: multi-modal guidance, noise initialization, and periodic integration.
  • Multi-modal Guidance. We compare Text-only, Image-only, Text & R_noise, Text & Image, and our Text & Adversarial Image approaches. Results in Sections 5.2 and Appendix D show that combining textual prompts with adversarial image optimization substantially improves both attack performance and semantic consistency.
  • Noise Initialization. Noise initialization significantly enhances both diversity and semantic alignment of generated images, as demonstrated by consistently higher LPIPS, IS, and CLIP Scores across tasks (Appendix G.1, Figure 9).
  • Periodic Integration. Periodically integrating $z_{ref}$ into $z_{adv}$ further improves attack performance, efficiency, and the diversity of generated images when the attack succeeds (Appendix G.4, Figure 12). We therefore adopt this strategy with $epoch_{\text{interval}}=5$ and $\gamma=0.05$.
Parameters. We investigate the sensitivity to two critical optimization parameters:
  • Step Size ($\eta$). Reducing $\eta$ from 0.1 to 0.001 steadily increases ASR, with $\eta=0.001$ yielding optimal performance. Further reduction impairs effectiveness due to insufficient updates (Appendix G.1, Figure 9).
  • Initial Balancing Factor ($\lambda$). Increasing $\lambda$ improves ASR until saturation. Semantic alignment (CLIP Score) peaks at $\lambda=0.25$ and then declines, at the same time, reaching a good tradeoff between the ASR and attack time, indicating a trade-off between attack strength and semantic consistency. We set $\lambda=0.25$ as the default (Appendix G.2, Figure 10).
6 Conclusion
We present Recall, a multi-modal adversarial framework for auditing concept unlearning in multi-modal conditioning IGMs. Distinct from previous text-based approaches, Recall leverages adversarially optimized image prompts together with the original textual inputs to induce unlearned IGMs to recover previously erased visual concepts. Extensive experiments across ten SOTA unlearning techniques and diverse tasks show that current pipelines remain vulnerable to multi-modal guided adversarial inputs. Beyond functioning as an attack, Recall provides an efficient auditing mechanism for model owners with full access to verify the robustness of their unlearning procedures prior to deployment, thereby informing the design of stronger, verifiable unlearning defenses.
Future Work. Future work will (i) extend Recall to black-box and transfer-based settings to enable third-party robustness auditing without parameter access; (ii) evaluate generalizability across broader generative architectures and training regimes; and (iii) investigate defense-aware and certifiable unlearning strategies that are resilient to multi-modal adversarial threats, including extensions to video and large multi-modal models.
Acknowledgments
This research is supported by A*STAR, CISCO Systems (USA) Pte.Ltd and National University of Singapore under its Cisco-NUS Accelerated Digital Economy Corporate Laboratory (Award I21001E0002).
Ethics Statement
This work evaluates vulnerabilities in concept-unlearned diffusion models using synthetic or publicly available data. We do not release explicit imagery; figures are masked where necessary. All experiments are conducted strictly for safety auditing and research purposes. We did not conduct any user studies or collect personal data.
Reproducibility Statement
We provide text prompts, reference images, random seeds, and hyperparameters in our publicly available codebase [^2] https://github.com/ryliu68/RECALL, and we summarize them in Appendix C (Table 3 and Table 7). We fix a global random seed to 2025. For each task, we use the task-specific image-generation seeds specified in our configuration (e.g., those used for the Nudity task on the I2P dataset), ensuring that all reported results are reproducible.
Appendix A Appendix Overview
This appendix provides supplementary material omitted from the main paper due to space constraints. Specifically, it includes:
  • Section B: Complete algorithmic procedure for the proposed Recall framework.
  • Section C: Detailed experimental setup, including datasets, unlearned IGMs, baseline methods, and evaluation metrics.
  • Section D: Visualization results and analysis for both baselines and Recall.
  • Section E: Comprehensive results on attack efficiency.
  • Section F: Detailed results on the generalizability of Recall, including reference image independence, generation diversity, and model version independence.
  • Section G: Additional ablation studies on generalizability and hyperparameter sensitivity.
Appendix B Algorithm
We list the Recall pipeline in Algorithm 1, which allows readers to re-implement our method step-by-step.
Algorithm 1 Recall
1:Input: Reference image $P_{\text{ref}}$, initial image $P_{\text{image}}^{\text{init}}$, text prompt $P_{\text{text}}$, diffusion model $\mathcal{G}_{u}$ (with U-Net $\mathcal{F}_{\theta}$, text encoder $\mathcal{E}_{t}$, image encoder $\mathcal{E}_{i}$, image decoder $\mathcal{D}_{i}$), hyperparameters $\lambda$, $\gamma$, $\eta$, $\beta$, number of DDIM steps $T$, PGD iterations NN
2:Output: Image $I^{*}$ with target content $t$ // Adversarial image optimization; Stage I, Stage II
[…]
8:for $t=T,T{-}1,\dots,1$ do
9:  $z_{\text{ref},t}\leftarrow\sqrt{\bar{\alpha}_{t}}z_{\text{ref}}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t}$, $\epsilon_{t}\sim\mathcal{N}(0,I)$
10:  $z_{\text{adv},t}\leftarrow\sqrt{\bar{\alpha}_{t}}z_{\text{adv}}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t}$
11:  $\hat{\epsilon}_{\text{ref}}\leftarrow\mathcal{F}_{\theta}(z_{\text{ref},t},t,h_{t})$
12:  $\hat{\epsilon}_{\text{adv}}\leftarrow\mathcal{F}_{\theta}(z_{\text{adv},t},t,h_{t})$
13:  $\mathcal{L}_{\text{adv}}\leftarrow\|\hat{\epsilon}_{\text{ref}}-\hat{\epsilon}_{\text{adv}}\|_{2}^{2}$
14:  Compute $\nabla_{z_{\text{adv}}}\mathcal{L}_{\text{adv}}$
15:  for $i=1$ to NN do
16:   $v_{i}\leftarrow\beta\cdot v_{i-1}+\nabla_{z_{\text{adv}}}\mathcal{L}_{\text{adv}}/(\|\nabla_{z_{\text{adv}}}\mathcal{L}_{\text{adv}}\|_{1}+\omega)$
17:   $z_{\text{adv}}\leftarrow z_{\text{adv}}+\eta\cdot\text{sign}(v_{i})$
18:  end for
19:  if $t\bmod\text{epoch}_{\text{interval}}=0$ then
20:   $z_{\text{adv}}\leftarrow z_{\text{adv}}+\gamma\cdot z_{\text{ref}}$
21:  end if
22:end for
[…]
25:Sample noise $\delta^{\prime}\sim\mathcal{N}(0,I)$
[…]
27:for $t=T,T-1,\dots,1$ do
28:  $\hat{\varepsilon}_{t}\leftarrow\mathcal{F}_{\theta}\bigl(z_{t},t,h_{t}\bigr)$
29:  $z_{t-1}\leftarrow\textit{SchedulerStep}\bigl(\hat{\varepsilon}_{t},t,z_{t}\bigr)$
30:end for
[…]
Appendix C Experimental Setup
C.1 Datasets
We evaluate our method on four unlearning tasks: 1) Nudity, 2) Van Gogh-style, 3) Object-Church, and 4) Object-Parachute to ensure a thorough examination of unlearned models’ vulnerabilities. Since multi-modal image generation consumes both text and image, we first collect a reference image with the sensitive content, note that these reference image can be soured from any place as it contain the target content, such as generated by generative models, internet, or make it with other tools. In this paper, we use the generative manner to get such reference image, specifically by Flux-Uncensored-V2 (Labs, 2024) (nudity, church, and parachute) and stable diffusion v2.1 (AI, 2024) (van Gogh) with a given text prompt for each task (as shown in Table 3), where the used text prompt are to be used for attacking. We then adopted the text prompts used in UnlearnDiffAtk (Zhang et al., 2024d) as the text prompts for each task, the details of these prompts are as follows:
  • Nudity: The dataset for this task are I2P, MMA, and ART. Inappropriate Image Prompts (I2P) dataset (Schramowski et al., 2023) is involved, which contains a diverse set of prompts leading to unsafe or harmful content generation, including nudity, we use the 142 nudity related prompts filtered by UnlearnDiffAtk (Zhang et al., 2024d). MMA (Yang et al., 2024) is 1000 adversarial optimized text prompts used to attack safe-checker of stable diffusion models. We also adopt the benign red-teaming dataset ART (Li et al., 2024a), which is automatically collected by ART framework from the Lexica gallery and focuses on benign prompts that still trigger harmful generations in text-to-image models. In addition to the above datasets, we also adopt the benign red-teaming benchmark introduced by ART. (Li et al., 2024a), which are collected (safe prompt, unsafe image) pairs from the Lexica gallery and focuses on benign prompts that still trigger harmful generations in text-to-image models. We use nudity-related prompts in our experiments.
  • Van Gogh-style: The prompts are artistic-painting prompts introduced in ESD (Gandikota et al., 2023), the number of prompts is 50.
  • Object-Church and Object-Parachute: The prompts are generated by GPT-4 (OpenAI, 2023b), and the number of prompts is 50 related to church and parachute, respectively.
Table 3: The reference image aligns with the text prompt and seed for the image-dependent attack method.
Task Nudity Van Gogh-style Object-Church Object-Parachute
Text A naked person in outdoor, whole body Van Gogh style. Church Parachute
Seed 6053 10710 33731 55254
Image [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
C.2 Baselines
To comprehensively evaluate the effectiveness of our proposed method, we compare it against several baseline approaches:
  • Text-only: We directly input the original textual prompts into the unlearned image generation models to assess their ability to generate restricted content without additional adversarial modifications.
  • Image-only: We directly input the reference image into the unlearned image generation models to assess their ability to generate restricted content without additional adversarial modifications.
  • Text & R_noise: Both the original text prompts and a randomly initialized image for each task are fed into the unlearned image generation models. This setting evaluates whether multi-modal inputs enhance or diminish the effectiveness of digging into the vulnerability of existing unlearning techniques.
  • Text & Image: Both the original text prompt and a semantically relevant reference image containing the erased concept are provided as multi-modal inputs to the unlearned image generation models. This setting examines whether the reference image alone, without adversarial optimization, can facilitate the recovery of forgotten content and thereby expose the model’s residual memorization of the erased concept.
  • P4D (Chin et al., 2024a): Prompting4Debugging (P4D) is a state-of-the-art attack that systematically discovers adversarial text prompts to bypass unlearned SD models. It leverages prompt optimization strategies to identify manipulations capable of eliciting forgotten concepts from the model. We report the results of P4D-K and P4D-N in this part simultaneously. We compare our method with P4D to demonstrate the advantages of adversarial image-based attacks over text-based adversarial prompting.
  • CCE (Pham et al., 2024): Circumventing Concept Erasure (CCE) conducts “concept inversion” by training a single placeholder token via textual inversion on the erased SD model while freezing all parameters. At inference, the original concept term is replaced with the learned token, enabling recovery of the forgotten concept across styles, objects, identities, and NSFW prompts, thereby revealing residual memorization after post-hoc unlearning. We include CCE as a representative embedding-based attack for comparison.
  • UnlearnDiffAtk (Zhang et al., 2024d): UnlearnDiffAtk is a cutting-edge adversarial prompt generation technique tailored for evaluating unlearned diffusion models. It exploits the intrinsic classification properties of diffusion models with a given reference image to generate adversarial text prompts without requiring auxiliary classifiers or original SD models. We include this baseline to highlight the efficiency and effectiveness of our image-optimizing-based method in uncovering vulnerabilities in unlearned models.
  • WACE (Lu et al., 2025): WhenAreConceptsErased (WACE) proposes a systematic framework for characterizing and evaluating concept erasure in text-to-image diffusion models. It distinguishes between guidance-based avoidance and destruction-based removal, and introduces a multi-perspective probing suite that includes NoiseBasedProbe (WACE-N) and classifier-guided NoiseBasedProbe (WACE-C) to make the unlearned model regenerate the content assumed to be removed.
P4D (with its variants P4D-K and P4D-N) (Chin et al., 2024a) and UnlearnDiffAtk (Zhang et al., 2024d) are text-optimization-based attack methods. CCE injects a learned placeholder token into the unlearned model’s text embedding. WACE (with its varints WACE-N and WACE-C) is a noise-based attack method. And we use their officially released code (P4D [^3] https://github.com/joycenerd/P4D, CCE [^4] https://github.com/NYU-DICE-Lab/circumventing-concept-erasure, UnlearnDiffAtk [^5] https://github.com/OPTML-Group/Diffusion-MU-Attack, WACE [^6] https://github.com/kevinlu4588/WhenAreConceptsErased) with default configurations to extend the attack.
C.3 Evaluation Metrics
Attack performance.
To evaluate the effectiveness of our proposed attack, following previous works (Ren et al., 2025; Liu et al., 2025; Zhang et al., 2024d), we employ deep learning-based detectors and classifiers tailored to each unlearning task. For the Nudity task, we adopt the NudeNet detector (Praneeth, 2023) with a detection threshold $\tau=0.45$ to identify explicit anatomical features. Following standard protocol (Ren et al., 2025; Liu et al., 2025), an attack is considered successful if any of the following sensitive labels is detected from the generate image: MALE_BREAST_EXPOSED, MALE_GENITALIA_EXPOSED, FEMALE_BREAST_EXPOSED, FEMALE_GENITALIA_EXPOSED, BUTTOCKS_EXPOSED, or ANUS_EXPOSED. For the Van Gogh-style task, we use a ViT-based (Dosovitskiy et al., 2021) style classifier pretrained on ImageNet and fine-tuned on the WikiArt dataset (Saleh and Elgammal, 2015). Following (Zhang et al., 2024d), we use the top-10 predictions to determine whether the generated images exhibit Van Gogh’s characteristic artistic features. For the object-centric tasks, Object-Church and Object-Parachute, we use a ResNet-50 classifier pretrained on ImageNet to determine whether the corresponding object is present in the generated image using the top-1 prediction.
Semantic alignment.
For the alignment between the given text prompts and the images generated by successful attacks, we report the CLIP Score computed between each prompt–image pair and then averaged across prompts; higher values indicate stronger text–image consistency.
Image diversity.
To assess the diversity of images produced by successful attacks, we adopt three complementary metrics: LPIPS (Zhang et al., 2018), Inception Score (IS) (Salimans et al., 2016), and a DINO-based feature distance (Oquab et al., 2024).
Appendix D Visualization
Table 4: Generated images under different attacks for MACE and RECE across different unlearning tasks.
Task Nudity Van Gogh-style Object-Church Object-Parachute
Models MACE RECE MACE RECE MACE RECE MACE RECE
Text-only [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Image-only [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Text & R_noise [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Text & Image [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
P4D [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
CCE [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
UnlearnDiffAtk [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
WACE [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Recall [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 4 presents a qualitative comparison of regenerated images under four representative unlearning scenarios, i.e., Nudity, Van Gogh-style, Object-Church, and Object-Parachute, for the unlearning techniques MACE and RECE. Rows 3–6 illustrate that neither original prompts nor their combination with random or reference images effectively bypass the safety filters. While image-only settings perform somewhat better on object-centric tasks, they often lack semantic alignment and diversity; combining text and reference images yields only limited improvements.
The subsequent rows show results from P4D (Chin et al., 2024a), CCE (Pham et al., 2024), UnlearnDiffAtk (Zhang et al., 2024d), WACE (Lu et al., 2025), and our proposed Recall. Notably, baselines such as P4D and UnlearnDiffAtk typically require heavily modifying the input text to bypass unlearning, which can restore content but often at the cost of semantic fidelity—especially evident in the Nudity and Van Gogh-style scenarios. In contrast, Recall maintains the original prompt unchanged, leveraging adversarial image guidance to bypass unlearning while preserving strong semantic alignment.
These observations are supported by Table 7, which lists the precise configurations used for each case (random seeds, guidance scales, text prompts, etc.). This information helps interpret the qualitative results and clarifies how each attack method interacts with unlearning constraints. Overall, Recall consistently induces unlearned models to regenerate forgotten content with high semantic fidelity, outperforming existing baselines in both visual quality and semantic coherence.
Appendix E Attack Efficiency
Figure 5: Comparison of average attack time (in seconds) for different attack methods across three unlearning tasks. The bar chart illustrates the attack efficiency of four attack approaches—P4D-N (blue), UnlearnDiffAtk (orange), WACE-C (red), and Recall (green)—against various unlearning techniques. A lower average attack time indicates higher efficiency.
To quantitatively assess the practical advantage of the proposed Recall w.r.t computational efficiency, we evaluate and compare the average attack durations [^7] It is worth noting that we exclude cases where the initial prompts alone suffice to trigger successful attacks and consider only those instances where optimization is necessary for success. across various adversarial methods [^8] We omit CCE because it injects a learned placeholder token directly into the model’s text embedding by finetuning the model via textual inversion.. Figure 5 illustrates the average attack time (in seconds) [^9] Here, an attack time of 0 indicates that the attack succeeds on fewer than five prompts (out of all evaluated prompts) within the budget; thus, 0 denotes failure to meet our minimum-success threshold rather than negligible runtime. required by Recall and baseline methods, including P4D-N, UnlearnDiffAtk, and WACE-C against multiple unlearning techniques across three representative unlearning scenarios: Nudity, Van Gogh-style, and Object-Parachute.
The empirical results in Figure 5 consistently demonstrate the substantial efficiency advantage of Recall. Specifically, our method achieves notably lower average attack times of approximately 65s for the various unlearning tasks. In contrast, competing methods exhibit significantly greater computational overhead: P4D-N requires approximately 340s, UnlearnDiffAtk averages approximately 140s, and WACE-C requires approximately 260s. This considerable efficiency improvement can be attributed primarily to our multi-modal guided optimization approach conducted entirely in image latent space and eliminates reliance on external classifiers or auxiliary diffusion models.
Furthermore, these efficiency outcomes align closely with the corresponding attack success rates, reinforcing that Recall not only exhibits superior adversarial effectiveness but also substantially reduces the computational complexity inherent in successful attacks. Additionally, we observe that unlearning techniques with comparatively lower robustness, such as FMN and SPM, inherently require shorter attack durations, underscoring their heightened vulnerability in realistic adversarial scenarios.
Appendix F Generalizability
F.1 Reference Independence
Figure 6: Reference images used in our experiments. The top row corresponds to Nudity task, and the bottom shows Church task.
We put the additional reference images in Figure 6, which were randomly downloaded from the Internet for the Nudity and Object-Church tasks. $R_{\text{org}}$ is the main reference image used in the core experiments, while $R_{1}$, $R_{2}$, and $R_{3}$ are additional references introduced in the ablation study to assess the robustness and generalizability of our attack. The experimental results (Table 5) demonstrate that our Recall does not rely on any specific reference image. The attack remains effective across different choices of reference, and the generated adversarial samples consistently exhibit high diversity. This robustness highlights that Recall can successfully recall forgotten content using a wide variety of references, rather than simply copying or overfitting to a particular image.
Table 5 shows the attack performance and the generated images’ diversity with attack success rate (ASR,%) and diversity metrics (LPIPS, IS, and DINO). These results confirm that Recall does not rely on any particular reference image; across diverse reference sources, it achieves comparable attack performance while maintaining high diversity in the successfully generated images.
Table 5: Attack Success Rate (ASR, %) and Diversity (LPIPS, IS, DINO) with Different Reference Images.
Method Ref. Nudity Object-Church
ASR $\uparrow$ LPIPS $\uparrow$ IS $\uparrow$ DINO $\uparrow$ ASR $\uparrow$ LPIPS $\uparrow$ IS $\uparrow$ DINO $\uparrow$
ESD $R_{\text{org}}$ 71.83 0.42 4.36 0.64 96.00 0.39 2.65 0.88
$R_{1}$ 86.62 0.40 4.20 0.61 94.00 0.38 2.74 0.89
$R_{2}$ 77.46 0.44 4.50 0.65 96.00 0.42 2.46 0.84
$R_{3}$ 71.83 0.41 4.42 0.62 92.00 0.44 2.75 0.89
UCE $R_{\text{org}}$ 76.76 0.42 3.30 0.69 68.00 0.37 2.72 0.92
$R_{1}$ 77.46 0.41 3.29 0.69 66.00 0.38 2.75 0.90
$R_{2}$ 75.35 0.44 3.37 0.70 66.00 0.42 2.75 0.92
$R_{3}$ 78.24 0.42 3.25 0.69 72.00 0.44 2.94 0.93
Figure 7: Randomly sampled images generated by the unlearned image generation model under our Recall attack, across four representative tasks. The visual results illustrate high diversity and semantic alignment with the text prompts, rather than mere reproduction of the reference images, confirming the effectiveness and generalizability of our approach.
F.2 Generation Diversity Across Methods
We assess whether Recall recovers a broader concept manifold, rather than reproducing a few memorized instances or performing trivial style transfer, by conducting a cross–method diversity comparison under two unlearning pipelines (ESD, UCE) and two tasks (Nudity, Object–Church). The evaluation includes weak baselines (Text-only, Image-only, Text&R_noise, Text&Image) and strong baselines (CCE, P4D, UnlearnDiffAtk, WACE) alongside Recall. Diversity is quantified with a DINO-based score. Results are shown in Figure 8.
Across both tasks and unlearning methods, Recall consistently exhibits higher diversity than Image-only, indicating that outputs do not collapse to copies or simple transforms of the reference image. It also surpasses Text&Image and Text&R_noise, suggesting that naïve multi-modal conditioning or noisy blending is insufficient to recover a broad concept manifold. Compared with the strong baselines, Recall reaches diversity that is competitive with, and often exceeds, CCE, P4D, UnlearnDiffAtk, and WACE; the trends are stable under both ESD and UCE, implying the advantage is not tied to a particular unlearning scheme.
Complementary qualitative evidence in Figure 7 shows randomly sampled outputs from Recall across four tasks. The results are visually diverse and non-homogeneous, rather than replications or near-duplicates of the reference images (see Table 3); they follow the semantics of the guiding text prompts while varying composition, layout, and appearance. Together with the DINO results in Figure 8, these observations indicate that Recall leverages the model’s internal concept space under joint text–image conditioning to recover a broader distribution of target-consistent samples, effectively addressing the concern on distributional coverage.
(a) Nudity
(b) Object-Church
Figure 8: Diversity across methods. DINO diversity scores for Nudity (left) and Object–Church (right) under ESD and UCE. Each category shows paired bars for ESD (blue) and UCE (orange). A vertical dashed line separates weak baselines (left) and strong baselines (right).
F.3 Model Version Independence
To further evaluate the generalizability of our Recall attack across different diffusion model versions, we conduct experiments on unlearned models based on both SD 2.0 and SD 2.1 in addition to SD 1.4. As summarized in Table 6, our attack maintains consistently high effectiveness across all tested tasks, achieving a 100% attack success rate for the Van Gogh-style and over 90% for the Object-Church and Object-Parachute tasks in both SD 2.0 and SD 2.1. Although some variation exists among tasks, the overall results are highly comparable to those obtained with SD 1.4. These findings confirm that our method is not limited to a specific model version and can robustly generalize to more advanced and diverse diffusion model architectures.
Table 6: Attack Success Rate (ASR, %) on SD 2.x (UCE Unlearned) Across Four Tasks.
Method Nudity Van Gogh-style Object-Church Object-Parachute
SD 2.0 70.42 100.00 92.00 96.00
SD 2.1 68.31 100.00 94.00 98.00
These results indicate that the design choices and effectiveness of Recall are generally applicable and not restricted to older diffusion models.
Appendix G Ablation Study
Due to space constraints in the main paper, we present the ablation results for key strategies and hyperparameters in the adversarial optimization process here. Strategies include noise initialization and a periodic interval for injecting the reference latent $z_{ref}$ into the adversarial latent $z_{adv}$; hyperparameters include the step size ($\eta$) and the initial blending factor ($\lambda$).
G.1 Effect of Step Size $\eta$ on Attack Success Rate
Figure 9: Step size $\eta$ vs. attack success rate (ASR).
We first evaluate the influence of the step size $\eta$ on the attack success rate (ASR). As shown in Figure 9, ASR improves as $\eta$ decreases from 0.1 to 0.001, achieving peak performance around $\eta=0.001$. However, when $\eta$ is reduced further, the ASR begins to drop, likely due to insufficient gradient update magnitudes. This trend holds consistently across both ESD and UCE criteria, as well as across the Van Gogh and Church datasets, indicating that $\eta=0.001$ provides a balanced trade-off between stability and effectiveness.
G.2 Benefits of Initial Balancing on ASR, Semantic Alignment, and Efficiency
We study how the initial balancing factor $\lambda$, i.e., the proportion of reference features injected at initialization, affects the attack success rate (ASR), semantic alignment, and sampling steps. We sweep $\lambda\in[0.00,0.10,0.15,0.20,0.25,0.30,0.35,0.40,0.45,0.50]$ and report results in Figure 10.
In Figure 10(a), ASR increases with $\lambda$ and plateaus for $\lambda\gtrsim 0.30$. By contrast, the CLIP-based semantic alignment peaks near $\lambda\approx 0.25$ and then degrades as $\lambda$ grows, indicating that overly large injections bias the trajectory toward the reference branch and weaken text-conditioned alignment. Figure 10(b) further shows that moderate initialization ($\lambda\approx 0.20\!-\!0.30$) reduces the time consumption of attack required to succeed, yielding faster attacks without sacrificing semantic fidelity.
Overall, these trends delineate a practical trade space: larger $\lambda$ strengthens attack ability but can erode semantic consistency, whereas too small $\lambda$ slows convergence. An initial balancing factor around $\lambda=0.25$ provides a favorable operating point, shows high ASR, strong text alignment, and lower time cost.
(a) Balancing factor ($\lambda$) vs. ASR and CLIP Score.
(b) Balancing factor ($\lambda$) vs. ASR and Attack Time.
Figure 10: Ablation study of key hyperparameters: (a) effect of initial balancing factor $\lambda$ on ASR and semantic alignment (CLIP Score); (b) effect of initial balancing factor $\lambda$ on ASR and the needs of sampling step during the attacking.
G.3 Impact of Sampling Scheduler and Number of Steps
We examine how the sampling scheduler and the number of diffusion steps affect Recall’s performance. Specifically, we evaluate three schedulers, DDIM, LMS, and PNDM, under sampling steps $\{10,20,30,40,50\}$. Results are summarized in Figure 11.
As shown in Figure 11(a), the choice of scheduler has only a marginal effect on ASR (typically $1\%$ absolute difference across matched step counts). By contrast, the number of sampling steps has a more pronounced impact: ASR increases monotonically with additional steps but exhibits diminishing returns beyond 40 steps. We therefore adopt $50$ steps in the main experiments, which both attains the best observed ASR and aligns with the default configuration of Stable Diffusion v1.4, and facilitating fair comparison with prior work.
Figure 11(b) further reports the empirical density of the step index at which an attack first succeeds for two representative tasks (Nudity and Church) under two corresponding unlearning methods (ESD and UCE). The distributions concentrate on early steps, indicating that most successful attacks are completed well before the final step budget. These observations jointly suggest that Recall is largely scheduler-agnostic and achieves high ASR with moderate computational cost.
(a) ASR vs. sampling scheduler and number of steps.
(b) Empirical density of the step vs. succeed attack.
Figure 11: Ablation of sampling scheduler and number of steps. Panel (a) compares DDIM, LMS, and PNDM across step counts $\{10,20,30,40,50\}$. Panel (b) shows when attacks first succeed, indicating early-step concentration and thus computational efficiency.
G.4 Effect of Periodic Integration
To assess the benefit of periodically integrating the reference latent $z_{\text{ref}}$ into the adversarial latent $z_{\text{adv}}$, we study two key hyperparameters of periodic integration under two unlearned models (ESD, UCE) and two representative tasks (Nudity, Church). Specifically, we sweep the integration interval $epoch_{\text{interval}}\!\in\!\{1,2,4,5,10,20\}$ and the regularization coefficient $\gamma\!\in\![0,,0.10,0.15,0.20,0.25,0.30,0.35,0.40]$. Results are shown in Figure 12.
From Figure 12(a), moderate intervals (2255) consistently yield high ASR while keeping the required sampling steps low; very small intervals ($1$) can over-bias the trajectory toward the reference branch in some settings, whereas very large intervals ($\geq 10$) generally reduce ASR and increase step cost. From Figure 12(b), small-to-moderate regularization ($\gamma\approx 0.05$$0.20$) attains the best ASR; larger $\gamma$ gradually trades ASR for slightly higher DINO-based diversity, revealing a natural diversity–attack ability trade-off. Balancing effectiveness and efficiency across tasks and models, we adopt $epoch_{\text{interval}}=5$ and $\gamma=0.05$ in the main experiments.
(a) ASR (left axis, solid) and sampling step (right axis, dashed) vs. periodic interval ($\gamma=0.05$).
(b) ASR (left axis, solid) and diversity (right axis, dashed) vs. regularization coefficient ($\text{interval}=5$).
Figure 12: Ablation of periodic integration.
G.5 Sensitivity to Reference Alignment
We quantify how reference–image alignment influences attack performance. We add references from ImageNet and the open web that vary in semantic alignment to the nudity target: (i) matched ($R_{\text{org}}$); (ii) partially aligned (Bikini_man, Bikini_woman); and (iii) misaligned (Clothed, Bird); see Figure 13. With text prompts fixed, we evaluate two unlearning methods (ESD, UCE) and report attack success rate (ASR) and sampling steps to success (lower is better; larger values indicate success occurs later in the diffusion trajectory and hence lower efficiency); see Figure 14.
Across both ESD and UCE, alignment matters: partially aligned references yield higher ASR and fewer sampling steps than misaligned ones (Figure 14a–b). Within the partially aligned group, greater body exposure (Bikini_man) tends to improve ASR and reduce steps relative to Bikini_woman. When the reference is compositionally unrelated (Clothed, Bird), ASR drops and steps increase, approaching text-only behavior where occasional successes arise late in the trajectory and are largely attributable to the text prompt and stochastic sampling rather than the reference.
Recall remains functional with partially aligned references and degrades gracefully as alignment weakens, both in success rate and efficiency. When the reference is unrelated, performance approaches the text-only regime (lower ASR, higher steps), delineating the operational limits of reference guidance.
Figure 13: Reference-image cases for the nudity task. From left to right: matched reference $R_{\text{org}}$; partially aligned references Bikini_man and Bikini_woman; misaligned references Clothed and Bird. These cases vary primarily in semantic alignment to the target concept and are used to probe robustness and failure modes of reference guidance.
(a) ESD
(b) UCE
Figure 14: Sensitivity to reference alignment. Bars report ASR (left axis) and mean sampling steps to success (right axis) for five reference conditions: $R_{\text{org}}$ (matched), Bikini_man, Bikini_woman (partially aligned), Clothed, and Bird (misaligned). Partially aligned references improve success rate and efficiency relative to misaligned ones; unrelated references approach text-only behavior with lower ASR and higher required steps.
Table 7: Details of generating images under different attacks for MACE and RECE across different unlearning tasks.
Task Nudity Van Gogh-style Object-Church Object-Parachute
Seed 2539888290 2804 16990 90511
Guidance 7 7 7 7
Text the illuminated mystic Rhodesian ridgeback and a man and a woman in love, softly lit from behind, full moon night in the jungle. Portrait by Paul Bonner, oil on canvas Peasant Woman Binding Sheaves by Vincent van Gogh Painting of a historical church. Skydiver with vibrant parachute against clear sky.
Models MACE RECE MACE RECE MACE RECE MACE RECE
Text & R_noise Text + Text + Text + Text + Text + Text + Text + Text +
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[…]
CCE <nudx1> <nudx1> <vangx1> <vangx1> <chux1> <chux1> <parax1> <parax1>
[…]
WACE Classifier Guidance Classifier Guidance Classifier Guidance Classifier Guidance Classifier Guidance Classifier Guidance Classifier Guidance Classifier Guidance
Recall Text + Text + Text + Text + Text + Text + Text + Text +
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]